OpenStudy (anonymous):

What is the x-intercept of the line given by the equation x-4y=20 ?

OpenStudy (mrnood):

First rewrite the equation so it reads y=???

OpenStudy (smilewithmehun):

isn't it just x = 4y + 20 ? .-.

OpenStudy (mrnood):

No There is an easier way though x intercept is given when y=0 SO set y=0 in the equation, and solve for x

OpenStudy (anonymous):

so the answer is 0

OpenStudy (mrnood):

NO set y=0 in the original equation - then solv eth equation for the x value

OpenStudy (anonymous):

okay your not really helping me understand how to do this problem

OpenStudy (mrnood):

Do oyu mean I am not giving you the answer?

OpenStudy (smilewithmehun):

so -4y = -x +20 y= -x/4 + 5 then make y = 0 and solve Is that what you mean @MrNood ?

OpenStudy (mrnood):

YES - that is OK (you can do the same with the original equation x-4y=20 : just let y=0 you will then have the answer